# Geometric Hardy inequalities on the Heisenberg groups via convexity

### v2

@@ Abstract.
We prove $L^{p}$-Hardy inequalities with distance to the boundary for domains in the Heisenberg group ${\mathbb{H}}^{n}$, $n\geq 1$. Our results are based on a certain geometric condition. This is first implemented for the Euclidean distance in certain non-convex domains. It is then implemented for the distance defined by the gauge quasi-norm related to the fundamental solution of the horizontal Laplacian when the domain is a half-space or a convex polytope. Finally it is implemented for the Carnot-Carathéodory distance on half-spaces and arbitrary bounded convex domains of ${\mathbb{H}}^{n}$. In all cases the constant $((p-1)/p)^{p}$ is obtained. In the more general context of a stratified Lie group of step two we study the superharmonicity and the weak $H$-concavity of the Euclidean distance to the boundary, thus obtaining an alternative proof for the $L^{2}$-Hardy inequality on convex domains.
@@ 1. Introduction
On the other hand, subelliptic Hardy inequalities have been studied for quite a long time and the work [GL90] of Garofallo and Lanconelli in the 90’s opened up the research in this direction. By subelliptic Hardy inequalities, we mean Hardy-type inequalities considered in the setting of homogeneous Lie groups, and in particular stratified groups. The systematic analysis of homogeneous Lie groups goes back to the seminal work [FS82] by Folland and Stein where the authors establish the corresponding “anisotropic” non-commutative harmonic analysis, see also [S93]. In view of their importance in the area of partial differential equations, stratified Lie groups, have been widely recognised as they play a key role in establishing subelliptic estimates for differential operators on general manifolds.
@@
where $d(x)$, $x\in\Omega$, stands for the Euclidean distance to $\partial\Omega$ and the constant $1/4$ is the best possible. Later on in [RSS20] Ruzhansky et al. proved that if $\Omega$ is a half-space in any stratified group then for any $p>1$ there holds
$\int_{\Omega}|\nabla_{H}u|^{p}{\rm d}x\geq\Big{(}\frac{p-1}{p}\Big{)}^{p}\int_{T}\frac{|\nabla_{H}d|^{p}}{d^{p}}|u|^{p}{\rm d}x\;,\qquad u\in{C^{\infty}_{c}}(\Omega),$
where $\nabla_{H}$ denotes the horizontal gradient on $H$. For other results in this direction see also [Rus18].
@@
In the case of the Euclidean distance on the Heisenberg group ${\mathbb{H}}^{n}$ we prove the following result which goes beyond the convexity condition. The precise value of the constant $\beta(p,n)$ is given in Proposition 3.
@@
Denoting by $d_{\rho}$ the Carnot-Carathéodory distance to the boundary, we obtain the following
@@
The appereance of the factor $|\nabla_{{\mathbb{H}}^{n}}d_{N}|$ in Theorem B – while a similar factor does not appear in Theorem C – should be seen in the light of a result of [MSC01] a special case of which states that $|\nabla_{{\mathbb{H}}^{n}}d_{\rho}|=1$ a.e. A more interesting difference between the above two homogeneous norms relates to nearest boundary points in the case of a subspace $\Pi\subset{\mathbb{H}}^{n}$. The correspondence
@@
where $s={\rm dist}(\xi,\partial\Pi)=d(\xi,\xi^{\prime})$ is different in the two cases. In the case of the quasi-norm (1) the above is a simple 1-1 correspondance, as in the Euclidean case, cf. Proposition 9. However in the case of the Carnot-Carathéodory distance the situation is entirely different; see Proposition 16.
@@
In our context the relevant notions of convexity of sets and functions are the ones introduced at the same time by Lu, Manfredi and Stroffolini in [LMS03] on the Heisenberg group and by Danielli, Garofallo and Nhieu in [DGN03] on any stratified group. These notions are the analogues of the corresponding ones in the abelian case $\mathbb{R}^{n}$ but with a twist; the condition refers to a convex combination of two elements $g,g^{\prime}$ for which, additionally, $g^{\prime}\in H_{g}$, i.e. $g^{\prime}$ lies in the horizontal plane passing through $g$; see Section 4.1 for the precise definitions. Exploring properties of the so-called weakly $H$-concave functions (see Definition 20) and using a result from [DGN03] we prove the following theorem, part (iii) of which is contained in [RSS20].
@@
${\rm(ii)}$ $\quad\Delta_{{\mathbb{H}}^{n}}d\leq 0\;\mbox{ in the distributional sense in }\Omega;$
@@
$\hskip 42.67912pt\int_{\Omega}|\nabla_{{\mathbb{H}}^{n}}u|^{2}{\rm d}g\geq\frac{1}{4}\int_{\Omega}\frac{|\nabla_{{\mathbb{H}}^{n}}d|^{2}}{d^{2}}u^{2}{\rm d}g\;,\qquad u\in{C^{\infty}_{c}}(\Omega),$
@@
Part (a) of the next theorem is essentially contained in [RSS20] but we include the short proof of it because of its central role in the present article.
@@
$d(g)-d(g^{\prime})$$\leq$ lim $\liminf\big{[}N(g^{-1}b_{n}^{\prime})-N(g^{\prime-1}b_{n}^{\prime})\big{]}$
$=$ lim $\liminf\big{[}N(g^{-1}g^{\prime}g^{\prime-1}b_{n}^{\prime})-N(g^{\prime-1}b_{n}^{\prime})\big{]}$
$\leq$ lim $\liminf\big{[}\beta N(g^{-1}g^{\prime})+N(g^{\prime-1}b_{n}^{\prime})-N(g^{\prime-1}b_{n}^{\prime})\big{]}$
@@
$d(g^{\prime})-d(g)\leq C\beta\,d_{N}g^{\prime},g)\,,$
@@ 3. Geometric Hardy inequalities on the Heisenberg group
In the section we consider geometric Hardy inequalities on the Heisenberg group $\mathbb{H}^{n}$. In the first part we consider the Euclidean distance and prove the validity of the Hardy inequality with best constant on certain torii under suitable assumptions on the radii. In the second part we study the geometric Hardy inequality for the pseudodistance induced by the quasi-norm (1).
@@
form the canonical basis basis of the first stratum and the associated horizontal gradient and horizontal Laplacian on ${\mathbb{H}}^{n}$ are given respectively by
@@
3.1. Hardy inequalities with respect to the Euclidean distance on a torus
We will see here that Theorem 1 can be applied in the case of the Euclidean distance on a torus and thus goes beyond the convexity assumption of [Lar16].
We first note that for any $u\in C^{2}({\mathbb{H}}^{n})$ we have
@@ 3.1. Hardy inequalities with respect to the Euclidean distance on a torus
and is smooth in $T$ except on the $(2n-1)$-dimensional ‘circle’
@@
(8) ×$\times\bigg{\{}\Big{[}(2n-1)(r-R)+4r^{3}\Big{]}(r-R)^{2}+\Big{[}r+(2n-1)(r-R)\Big{]}t^{2}\bigg{\}}.$
@@
Case $21<p<$ 2. In this case we choose
@@
Moreover if we define $p_{0}=\sqrt{9n^{2}-8n+2}-3(n-1)$ then 3/$23/2<p_{0}<$ 2 and for $21<p<$ 2 we have
@@
It is evident from the argument in the above proof that if $\Omega\subset\mathbb{H}^{n}$ is a domain with $C^{2}$ boundary then the corresponding Hardy constant connot be larger than $((p-1)/p)^{p}$, provided there exists a point $\xi_{0}\in\partial\Omega$ such that $\nabla_{\mathbb{H}^{n}}d(\xi_{0})\neq 0$. This, of course, is very generic. We do not pursue this any further here, but we make two comments.
1. Assume that $\Omega$ possesses cylindrical symmetry so that $d=d(r,t)$. Suppose that at a point $\xi_{0}\in\partial\Omega$ we have $\nabla_{\mathbb{H}^{n}}d(\xi_{0})=0$, that is
$d_{x_{i}}+2y_{i}d_{t}=0\;,\qquad d_{y_{i}}-2x_{i}d_{t}=0\;,\qquad i=1,\ldots,n.$
Multiplying by $x_{i}$ and $y_{i}$ respectively and adding we obtain
$\sum_{i=1}^{n}(x_{i}d_{x_{i}}+y_{i}d_{y_{i}})=0,\qquad\mbox{ at the point }\xi_{0}.$
Using cylindrical coordinates $(r,\omega,t)$ we then obtain that $d_{r}(\xi_{0})=0$. By cylindrical symmetry we have
$d_{r}^{2}+d_{t}^{2}=|\nabla d|^{2}=1\;,\qquad\mbox{ in }T,$
hence $d_{t}(\xi_{0})=1$ and therefore the tangent hyperplane at $\xi_{0}$ must be parallel to the hyperplane $\{t=0\}$. Hence generically we have $\nabla_{\mathbb{H}^{n}}d\neq 0$ on $\partial\Omega$
2. Let $\Omega$ be a domain which does not necessarily possess some kind of symmetry. If $\xi_{0}\in\partial\Omega$ is a boundary point with $\nabla_{\mathbb{H}^{n}}d(\xi_{0})=0$, then we have
@@ 3.2. Hardy inequalities with respect to the gauge pseudodistance
$d_{N}(\xi)=d_{N}(r,t)={\left\{\begin{array}[]{ll}{\big{(}2r^{4}s^{2}-3tr^{2}s+t^{2}\big{)}^{\frac{1}{4}},}&{r>0,\;t>0,}\\ {t^{\frac{1}{2}},}&{r=0,\;t>0,}\end{array}\right.}$
@@
For simplicity we will write $d$ instead of $d_{N}$. By group action (see also [Lar16, p340]) it is enough to consider the case $\Pi=\Pi_{0}$. Also, it is preferable to work with the function
$G(r,t)=d(r,t)^{4}$
instead of $d(r,t)$. To compute the various derivatives of $G(r,t)$ we recall from Lemma 16 that
@@
Setting $A=|\nabla_{\mathbb{H}^{n}}d|^{2}$ we have, cf. (9),
(25) $\Delta_{p,\mathbb{H}^{n}}d=A^{\frac{p-4}{2}}\Big{(}A\,\Delta_{\mathbb{H}^{n}}d+\frac{p-2}{2}\big{(}d_{r}A_{r}+4r^{2}d_{t}A_{t}\big{)}\Big{)}.$
@@
$d_{r}=\frac{1}{4}G^{-\frac{3}{4}}G_{r},\qquad d_{t}=\frac{1}{4}G^{-\frac{3}{4}}G_{t},\qquad d_{rr}=-\frac{3}{16}G^{-\frac{7}{4}}G_{r}^{2}+\frac{1}{4}G^{-\frac{3}{4}}G_{rr}$
@@
$d_{tt}=-\frac{3}{16}G^{-\frac{7}{4}}G_{t}^{2}+\frac{1}{4}G^{-\frac{3}{4}}G_{tt},\qquad d_{rt}=-\frac{3}{16}G^{-\frac{7}{4}}G_{r}G_{t}+\frac{1}{4}G^{-\frac{3}{4}}G_{rt}.$
@@
$A=d_{r}^{2}+4r^{2}d_{t}^{2}=\frac{1}{16}G^{-\frac{3}{2}}(G_{r}^{2}+4r^{2}G_{t}^{2})=G^{-\frac{3}{2}}r^{6}s^{6}(s^{2}+1),$
$A_{r}$ $=2d_{r}d_{rr}+8rd_{t}^{2}+8r^{2}d_{t}d_{rt}$
@@
$\Delta_{{\mathbb{H}}^{n}}d$ $=d_{rr}+\frac{2n-1}{r}d_{r}+4r^{2}d_{tt}$
@@
$\Delta_{p,\mathbb{H}^{n}}d=-G^{-\frac{3p+1}{4}}\,\frac{r^{3p}s^{3p+2}(s^{2}+1)^{\frac{p}{2}}\big{(}(6n+p-4)s^{2}+4n+p-5\big{)}}{3s^{2}+2}\leq 0,$
and the desired inequality has been proved pointwise in $\Pi_{0}\setminus\{r=0\}$ (where $d_{N}$ is smooth). To complete the proof we argue as in the proof of Theorem 4, using in particular functions $\psi_{\epsilon}$, $\epsilon>0$, as in that proof. Part (i) is also used at this point since the local boundedness of $|\nabla_{H}d|$ is required when letting $\epsilon\to 0$. ∎
@@
$\frac{1}{d^{6}(r,t)}=\frac{64r^{6}}{t^{6}}+O(\frac{1}{t^{4}})$
@@
$d_{t}^{2}=\frac{1}{4r^{2}}+O(t^{2})\;,\qquad d_{r}^{2}=\frac{t^{2}}{4r^{4}}+O(t^{4}).$
@@
For simplicity we write $d$ instead of $d_{N}$. Let $E_{1},\ldots,E_{m}$ denote the sides of $\Omega$. We define
@@
(28) $d(\xi)={\rm dist}(\xi,\partial\Pi_{k})=:d_{k}(\xi)\;,\qquad\mbox{ for all }\xi\in A_{k}.$
@@
(29) $\int_{\Omega}\nabla_{H}d\cdot\nabla_{H}\phi\,{\rm d}\xi=\sum_{k=1}^{m}\int_{A_{k}}\nabla_{H}d_{k}\cdot\nabla_{H}\phi\,{\rm d}\xi\geq 0.$
@@
We cannot directly apply integration by part to each of the integrals in the RHS of (29) due to the fact that, as is seen from Lemma 5, there is a halfline $L_{k}\subset\Pi_{k}$ on which each function $d_{k}$ is not differentiable (the halfline $L_{k}$ is the image of the halfline $\{(0,0,t):t>0\}$ under the group action that maps $\Pi_{0}$ onto $\Pi_{k}$). We define the set

### v3

@@ Abstract.
We prove $L^{p}$-Hardy inequalities with distance to the boundary for domains in the Heisenberg group ${\mathbb{H}}^{n}$, $n\geq 1$. Our results are based on a certain geometric condition. This is first implemented for the Euclidean distance in certain non-convex domains. It is then implemented for the distance defined by the gauge quasi-norm related to the fundamental solution of the horizontal Laplacian when the domain is a half-space or a convex polytope. Finally it is implemented for the Carnot-Carathéodory distance on half-spaces and arbitrary bounded convex domains of ${\mathbb{H}}^{n}$. In all cases the constant $((p-1)/p)^{p}$ is obtained. In the more general context of a stratified Lie group of step two we study the $p$-superharmonicity and the weak $H$-concavity of the Euclidean distance to the boundary, thus obtaining an $L^{p}$-Hardy inequality on bounded convex domains which generalises previous results.
@@ 1. Introduction
On the other hand, subelliptic Hardy inequalities have been studied for quite a long time and the work [GL90] of Garofalo and Lanconelli in the 90’s opened up the research in this direction. By subelliptic Hardy inequalities, we mean Hardy-type inequalities considered in the setting of homogeneous Lie groups, and in particular stratified groups. The systematic analysis of homogeneous Lie groups goes back to the seminal work [FS82] by Folland and Stein where the authors establish the corresponding “anisotropic” non-commutative harmonic analysis, see also [S93]. In view of their importance in the area of partial differential equations, stratified Lie groups, have been widely recognised as they play a key role in establishing subelliptic estimates for differential operators on general manifolds.
@@
where $d(x)$, $x\in\Omega$, stands for the Euclidean distance to $\partial\Omega$ and the constant $1/4$ is the best possible. Later on in [RSS20b] Ruzhansky et al. proved that if $\Omega$ is a half-space in $\mathbb{H}^{n}$ and $p>1$ then
$\int_{\Omega}|\nabla_{\mathbb{H}^{n}}u|^{p}{\rm d}x\geq\Big(\frac{p-1}{p}\Big)^{p}\int_{\Omega}\frac{|\nabla_{\mathbb{H}^{n}}d|^{p}}{d^{p}}|u|^{p}{\rm d}x\;,\qquad u\in{C^{\infty}_{c}}(\Omega).$
A little earlier the same authors had showed that if $\Omega$ is a convex domain in any stratified group then
$\int_{\Omega}|\nabla_{H}u|^{2}\,{\rm d}x\geq\frac{1}{4}\int_{\Omega}\frac{|\nabla_{H}d|^{2}}{d^{2}}u^{2}\,{\rm d}x\,$,
where $\nabla_{H}$ denotes the horizontal gradient [RSS20a]. For other results in this direction see also [Rus18]; see also [PRS19] for an alternative approach based on a sub-Riemannian version of the Santaló formula.
@@
Let us first consider some known results about Hardy inequalities in Euclidean domains. Assume that $U\subset\mathbb{R}^{d}$ is a bounded domain with $C^{2}$ boundary and let $p>1$.
One approach to study the Hardy inequality in $U$ is by using the fact that if we consider functions supported in a small neighbourhood of a boundary point, then the Hardy constant is $((p-1)/p)^{p}$. However, if we try to pass from this local inequality to a global one on $U$, then one ends up with the inequality
$\int_{U}|\nabla u|^{p}{\rm d}x+c\int_{U}|u|^{p}{\rm d}x\geq\Big(\frac{p-1}{p}\Big)^{p}\int_{U}\frac{|u|^{p}}{d^{p}}{\rm d}x\;,\qquad u\in{C^{\infty}_{c}}(U).$
for some $c>0$ depending on $U$. It is not possible to avoid the $L^{p}$ term unless some geometric assumption is made on $U$.
Indeed, for $p=2$, [Dav95] Davies constructed certain planar sectors with Hardy constant smaller that $1/4$. By approximating such a sector by an increasing sequence of bounded smooth domains one concludes the existence of bounded smooth domains with Hardy constant smaller than $1/4$. In the same article and for dimension $d\geq 3$, Davies constructed conical domains with Hardy constant arbitrarily small; again, this shows the existence of bounded smooth domains with the same property. Concerning the general case $p>1$, elemantary considarations show that when the domain is the complement of a ball in $\mathbb{R}^{d}$ and $p>(d+1)/2$ then the (global) Hardy constant is smaller than $((p-1)/p)^{p}$. A simple argument then shows the existence of tori for which the same holds true.
It is a central question to find conditions on the domain $U$ under which the Hardy constant is equal to $((p-1)/p)^{p}$.
In the context of the Heisenberg group $\mathbb{H}^{n}$ one similarly expects that in order to obtain the Hardy constant $((p-1)/p)^{p}$ some geometric condition is necessary on the domain. Indeed, Proposition 3 below uses a lifting argument to show that each of the above examples for Euclidean domains induces a corresponding example in $\mathbb{H}^{n}$ for the Hardy inequality with Euclidean distance. The main idea here is to lift the Euclidean domain $U\subset\mathbb{R}^{n}$ to the domain $U\times\mathbb{R}\subset{\mathbb{H}}^{n}$ in the Heisenberg group.
The condition introduced in this article, namely the $p$-horizontal superharmonicity of the distance function, turns out to be a natural condition. For the case of the Euclidean distance in a Carnot group of step two it applies to any convex domain for any $p>1$, see Section 4. But it is also valid to domains that are not necessarily convex as is seen in Section 2. Moreover it is applied, as already mentioned, to the Carnot-Carathéodory distance and the gauge pseudodistance.
We now present the main results of this article. The torus $T\subset\mathbb{H}^{n}$ defined in the next theorem is not a convex domain but satisfies the geometric assumption $\Delta_{p,\mathbb{H}^{n}}d\leq 0$ for any $p>1$. The precise value of the constant $\beta(p,n)$ is given in Proposition 4.
@@
$\Delta_{p,\mathbb{H}^{n}}\,d\leq 0\,\quad\text{in the weak sense in}\,\,T\,$,
and
@@
Denoting by $d_{\rho}$ the Carnot-Carathéodory distance (control distance) to the boundary, we obtain the following
@@
The appearance of the factor $|\nabla_{{\mathbb{H}}^{n}}d_{N}|$ in Theorem B – while a similar factor does not appear in Theorem C – should be seen in the light of a result of [MSC01] a special case of which states that $|\nabla_{{\mathbb{H}}^{n}}d_{\rho}|=1$ a.e. A more interesting difference between the above two homogeneous norms relates to nearest boundary points in the case of a subspace $\Pi\subset{\mathbb{H}}^{n}$. The correspondence
@@
where $s={\rm dist}(\xi,\partial\Pi)=d(\xi,\xi^{\prime})$ is different in the two cases. In the case of the quasi-norm (1) the above is a simple 1-1 correspondence, as in the Euclidean case, cf. Proposition 10. However in the case of the Carnot-Carathéodory distance the situation is entirely different; see Proposition 17.
@@
In our context the relevant notions of convexity of sets and functions are the ones introduced at the same time by Lu, Manfredi and Stroffolini in [LMS03] on the Heisenberg group and by Danielli, Garofalo and Nhieu in [DGN03] on any stratified group. These notions are the analogues of the corresponding ones in the abelian case $\mathbb{R}^{n}$ but with a twist; the condition refers to a convex combination of two elements $g,g^{\prime}$ for which, additionally, $g^{\prime}\in H_{g}$, i.e. $g^{\prime}$ lies in the horizontal plane passing through $g$; see Section 4.1 for the precise definitions. Exploring properties of the so-called weakly $H$-concave functions (see Definition 21) and using a result from [DGN03] we prove the following theorem, part (iii) of which generalizes [RSS20b, Corollary 2.2].
@@
${\rm(ii)}$ $\quad\Delta_{p,H}d\leq 0\;\mbox{ in the distributional sense in }\Omega;$
@@
$\hskip 42.67912pt\int_{\Omega}|\nabla_{H}u|^{p}{\rm d}g\geq\Big(\frac{p-1}{p}\Big)^{p}\int_{\Omega}\frac{|\nabla_{H}d|^{p}}{d^{p}}|u|^{p}{\rm d}g\;,\qquad u\in{C^{\infty}_{c}}(\Omega),$
@@
Part (a) of the next theorem is essentially contained in [RSS20b] but we include the short proof of it because of its central role in the present article.
@@
$d(g)-d(g^{\prime})$ $\leq$ $\liminf\big[N(g^{-1}b_{n}^{\prime})-N(g^{\prime-1}b_{n}^{\prime})\big]$
$=$ $\liminf\big[N(g^{-1}g^{\prime}g^{\prime-1}b_{n}^{\prime})-N(g^{\prime-1}b_{n}^{\prime})\big]$
$\leq$ $\liminf\big[\beta N(g^{-1}g^{\prime})+N(g^{\prime-1}b_{n}^{\prime})-N(g^{\prime-1}b_{n}^{\prime})\big]$
@@
$d(g^{\prime})-d(g)\leq C\beta\,d_{N}(g^{\prime},g)\,$,
@@ 3. Geometric Hardy inequalities on the Heisenberg group
In the section we consider geometric Hardy inequalities on the Heisenberg group $\mathbb{H}^{n}$. In the first part we consider the Euclidean distance and prove the validity of the Hardy inequality with best constant on certain tori under suitable assumptions on the radii. In the second part we study the geometric Hardy inequality for the pseudodistance induced by the quasi-norm (1).
@@
form the canonical basis of the first stratum and the associated horizontal gradient and horizontal Laplacian on ${\mathbb{H}}^{n}$ are given respectively by
@@
3.1. Hardy inequalities with respect to the Euclidean distance: remarks and counterexamples
In this section we consider geometric Hardy inequalities in domains $\Omega\subset\mathbb{H}^{n}$ with respect to the Euclidean distance.
The next proposition, when combined with the examples of Euclidean domains for which the Hardy constant is smaller than $((p-1)/p)^{p}$, shows the one cannot expect that a bounded $C^{2}$ domain in $\mathbb{H}^{n}$ has constant $((p-1)/p)^{p}$ unless some geometric condition is imposed.
Proposition 3.
Let $U\subset\mathbb{R}^{2n}$ be a domain. The Hardy constant of $\Omega=U\times\mathbb{R}$ as a domain in $\mathbb{H}^{n}$ cannot exceed the Hardy constant of $U$.
Proof.
Let $v\in C^{\infty}_{c}(U)$ and $\psi\in{C^{\infty}_{c}}(\mathbb{R})$ be given. We define
$u(x,y,t)=v(x,y)\psi(t)\;,\qquad(x,y)\in U,\;t\in\mathbb{R},$
A simple computation gives that for any $\epsilon>0$,
$|\nabla_{\mathbb{H}^{n}}u|^{2}$ $=|\nabla v|^{2}\psi^{2}+4r^{2}v^{2}(\psi^{\prime})^{2}+4v\psi\psi^{\prime}\sum_{k=1}^{n}(y_{k}v_{x_{k}}-x_{k}v_{y_{k}})$
$\leq(1+\epsilon)|\nabla v|^{2}\psi^{2}+4(1+\frac{1}{\epsilon})r^{2}v^{2}(\psi^{\prime})^{2}.$
Combining this with the elementary inequality
$|a+b|^{q}\leq(1+\epsilon)|a|^{q}+k_{\epsilon}|b|^{q}\;,\quad\epsilon>0,$
we obtain that for any $\epsilon>0$ there exists $c_{\epsilon}>0$ such that
$|\nabla_{\mathbb{H}^{n}}u|^{p}\leq(1+\epsilon)|\nabla v|^{p}|\psi|^{p}+c_{\epsilon}r^{p}|v|^{p}|\psi^{\prime}|^{p}\;,\quad\mbox{ in }U^{*}.$
Since $d_{\Omega}(x,y,t)=d_{U}(x,y)$, we conclude that
$\frac{\int_{\Omega}|\nabla_{\mathbb{H}^{n}}u|^{p}{\rm d}\xi}{\int_{\Omega}\frac{|u|^{p}}{d^{p}}{\rm d}\xi}\leq(1+\epsilon)\frac{\int_{U}|\nabla v|^{p}dx\,dy}{{\int_{U}\frac{|v|^{p}}{d^{p}}dx\,dy}}+c_{\epsilon}\frac{\int_{U}r^{p}|v|^{p}dx\,dy}{\int_{U}|v|^{p}dx\,dy}\cdot\frac{\int_{-\infty}^{\infty}|\psi^{\prime}|^{p}dt\,dy}{\int_{-\infty}^{\infty}|\psi|^{p}dt}.$
The result follows by noting that the quotient involving $\psi$ can be made arbitrarily small. ∎
We shall now consider the case of the torus $T$ which is not a convex domain and has been discussed in Theorem A. For $u\in C^{2}({\mathbb{H}}^{n})$ we have
@@ 3.1. Hardy inequalities with respect to the Euclidean distance: remarks and counterexamples
and is smooth in $T$ except on the (2n1)(2n-1)-dimensional ‘circle’
@@
(8) $\times\bigg\{\Big[(2n-1)(r-R)+4r^{3}\Big](r-R)^{2}+\Big[r+(2n-1)(r-R)\Big]t^{2}\bigg\}.$
@@
Case $1<p<2$. In this case we choose
@@
Moreover if we define $p_{0}=\sqrt{9n^{2}-8n+2}-3(n-1)$ then $3/2<p_{0}<2$ and for $1<p<2$ we have
@@
It is evident from the argument in the above proof that if $\Omega\subset\mathbb{H}^{n}$ is a domain with $C^{2}$ boundary then the corresponding Hardy constant cannot be larger than $((p-1)/p)^{p}$, provided there exists a point $\xi_{0}\in\partial\Omega$ such that $\nabla_{\mathbb{H}^{n}}d(\xi_{0})\neq 0$. This is very generic. We refer to [M06] and references therein for results concerning the negligibility of the set of such points in a much wider context.
In the case of the Heisenberg group $\mathbb{H}^{n}$ we can give a direct proof of this negligibility. Indeed, if $\xi_{0}\in\partial\Omega$ is a boundary point with $\nabla_{\mathbb{H}^{n}}d(\xi_{0})=0$, then we have
@@ 3.2. Hardy inequalities with respect to the gauge pseudodistance
(16) $d_{N}(\xi)=d_{N}(r,t)={\left\{\begin{array}[]{ll}{\big(2r^{4}s^{2}-3tr^{2}s+t^{2}\big)^{\frac{1}{4}},}&{r>0,\;t>0,}\\ {t^{\frac{1}{2}},}&{r=0,\;t>0,}\end{array}\right.}$
@@
By group action (see also [Lar16, p340]) it is enough to consider the case $\Pi=\Pi_{0}$. Also, it is preferable to work with the function
$G(r,t)=d_{N}(r,t)^{4}$
instead of $d_{N}(r,t)$. To compute the various derivatives of $G(r,t)$ we recall from Lemma 17 that
@@
Setting $A=|\nabla_{\mathbb{H}^{n}}d_{N}|^{2}$ we have, cf. (9),
(26) $\Delta_{p,\mathbb{H}^{n}}d_{N}=A^{\frac{p-4}{2}}\Big(A\,\Delta_{\mathbb{H}^{n}}d_{N}+\frac{p-2}{2}\big(d_{N,r}A_{r}+4r^{2}d_{N,t}A_{t}\big)\Big).$
@@
$d_{N,r}=\frac{1}{4}G^{-\frac{3}{4}}G_{r},\qquad d_{N,t}=\frac{1}{4}G^{-\frac{3}{4}}G_{t},\qquad d_{N,rr}=-\frac{3}{16}G^{-\frac{7}{4}}G_{r}^{2}+\frac{1}{4}G^{-\frac{3}{4}}G_{rr}$
@@
$d_{N,tt}=-\frac{3}{16}G^{-\frac{7}{4}}G_{t}^{2}+\frac{1}{4}G^{-\frac{3}{4}}G_{tt},\qquad d_{N,rt}=-\frac{3}{16}G^{-\frac{7}{4}}G_{r}G_{t}+\frac{1}{4}G^{-\frac{3}{4}}G_{rt}.$
@@
$A=d_{N,r}^{2}+4r^{2}d_{N,t}^{2}=\frac{1}{16}G^{-\frac{3}{2}}(G_{r}^{2}+4r^{2}G_{t}^{2})=G^{-\frac{3}{2}}r^{6}s^{6}(s^{2}+1),$
$A_{r}$ $=2d_{N,r}d_{N,rr}+8rd_{t}^{2}+8r^{2}d_{N,t}d_{N,rt}$
@@
$\Delta_{{\mathbb{H}}^{n}}d_{N}$ $=d_{N,rr}+\frac{2n-1}{r}d_{N,r}+4r^{2}d_{N,tt}$
@@
$\Delta_{p,\mathbb{H}^{n}}d_{N}=-G^{-\frac{3p+1}{4}}\,\frac{r^{3p}s^{3p+2}(s^{2}+1)^{\frac{p}{2}}\big((6n+p-4)s^{2}+4n+p-5\big)}{3s^{2}+2}\leq 0,$
and the desired inequality has been proved pointwise in $\Pi_{0}\setminus\{r=0\}$ (where $d_{N}$ is smooth). To complete the proof we argue as in the proof of Theorem 5, using in particular functions $\psi_{\epsilon}$, $\epsilon>0$, as in that proof. Part (i) is also used at this point since the local boundedness of $|\nabla_{H}d_{N}|$ is required when letting $\epsilon\to 0$. ∎
@@
$\frac{1}{d_{N}^{6}(r,t)}=\frac{64r^{6}}{t^{6}}+O(\frac{1}{t^{4}})$
@@
$d_{N,t}^{2}=\frac{1}{4r^{2}}+O(t^{2})\;,\qquad d_{N,r}^{2}=\frac{t^{2}}{4r^{4}}+O(t^{4}).$
@@
Remark 3.
We note for later use that we can use (17) to eliminate $t$ from the expressions (29), (29) and (30) for $d_{N}$, $d_{N,t}$ and $d_{N,r}$ respectively. We obtain
$d_{N}=rs(s^{2}+1)^{1/4},\qquad d_{N,r}=\frac{r^{2}s^{3}}{2d^{3}},\qquad d_{N,t}=\frac{r^{2}s^{3}}{2d^{3}}.$
@@
Let $E_{1},\ldots,E_{m}$ denote the sides of $\Omega$. We define
@@
(31) $d_{N}(\xi)={\rm dist}_{N}(\xi,\partial\Pi_{k})=:d_{N,k}(\xi)\;,\qquad\mbox{ for all }\xi\in A_{k}.$
@@
(32) $\int_{\Omega}\nabla_{H}d_{N}\cdot\nabla_{H}\phi\,{\rm d}\xi=\sum_{k=1}^{m}\int_{A_{k}}\nabla_{H}d_{N,k}\cdot\nabla_{H}\phi\,{\rm d}\xi\geq 0.$
@@
We cannot directly apply integration by part to each of the integrals in the RHS of (32) due to the fact that, as is seen from Lemma 6, there is a halfline $L_{k}\subset\Pi_{k}$ on which the function $d_{N,k}$ is not differentiable (the halfline $L_{k}$ is the image of the halfline $\{(0,0,t):t>0\}$ under the group action that maps $\Pi_{0}$ onto $\Pi_{k}$). We define the set